I need to now how to determine if an equation's line is parallel, perpendicular or neither. I have to do it without he use of a graph. I know how to determine if it is parallel because the slopes are the same and I have some understanding that perpendicular lines having opposite reciprocals. My question is, how do I tell the difference between a reciprocal and an opposite reciprocal.
Answer by stanbon(75887)

The opposite reciprocal is the negative of the reciprocal.
-----
For two lines to be perpendicular the product of the slopes must be -1.
m1*m2 = -1 so m1(-1/m1) = -1
So m2 = -1/m1
------------------
If one of the slopes is 2 the other must be -1/2
---------------
Cheers,
Stan H.